OpenStudy (anonymous):

How do I isolate "y" from the fractions, in order to put in slope intercept form 1/2 x - 2/3y=1

OpenStudy (anonymous):

Multiply out by 6

OpenStudy (anonymous):

Multiply by 3/2 to get 3/4 x -y = 3/2 so y= 3/4 x -3/2